OpenStudy (anonymous):

A segment has endpoints A (-1, 1) and B (8, 4) . Find the coordinates of the three points that divide the segment into four equal parts. (Named from A to B, X Y Z) (Use fraction format.) X = (, ) Y = (, ) Z = (, )

hero (hero):

I think I can help you with this.

OpenStudy (anonymous):

ok

hero (hero):

I know that Y = (9/2, 3/2)

hero (hero):

X = 5/4, 7/4

hero (hero):

sorry this is taking so long

hero (hero):

Z = 25/4, 11/4

hero (hero):

These were not easy to figure out

hero (hero):

Make sure you input these in the correct order: X = (5/4, 7/4) Y = (9/2, 3/2) Z = (25/4, 11/4)

hero (hero):

Johny, let me know if your computer likes the solution or not.

OpenStudy (anonymous):

it says partial credit so i think some of its ronq

hero (hero):

Oh well

OpenStudy (anonymous):

lol